### v1

# Parameter-Efficient Augment Plugin for Class-Incremental Learning

Abstract
Existing class-incremental learning (CIL) approaches based on replay or knowledge distillation are often constrained by forgetting or the stability-plasticity dilemma. Some expansion-based approaches could achieve higher accuracy. However, they always require significant parameter increases. In this paper, we propose a plugin extension paradigm termed the Deployment of extra LoRA Components (DLC) for non-pre-trained CIL scenarios. We treat the feature extractor trained through replay or distillation as a base model with rich knowledge. For each task, we use Low-Rank Adaptation (LoRA) to inject task-specific residuals into the base model’s deep layers. During inference, representations with task-specific residuals are aggregated to produce classification predictions. To mitigate interference from non-target LoRA plugins, we introduce a lightweight weighting unit. This unit learns to assign importance scores to different LoRA-tuned representations. Like downloadable contents in software, our method serves as a plug-and-play enhancement that efficiently extends the base methods. Remarkably, on the large-scale ImageNet‑100, with merely 4% of the parameters of a standard ResNet-18, our DLC model achieves a significant 8% improvement in accuracy, demonstrating exceptional efficiency. Moreover, it could surpass state-of-the-art methods under the fixed memory budget.
1 Introduction
In recent years, deep learning has achieved remarkable success across a variety of domains [jaimes2007multimodal, ye2019learning, chen2021large, chen2022learning, yang2024entaugment]. However, most models rely on a key assumption that the data distribution and learned classes remain fixed. In real-world scenarios, data distributions evolve continuously and new classes emerge over time [gomes2017survey], making static training paradigms impractical. Moreover, learning each new task typically relies on introducing additional task-specific sub-models, leading to inefficient model growth over time. Class-Incremental Learning (CIL) [zhou2024class, liangloss, yang2025supervised] thus plays a crucial role in enabling models to continually learn from non-stationary data streams while retaining prior knowledge.
Existing CIL methods  [zhou2024class] can be broadly categorized into rehearsal-based, restriction-based, and expansion-based approaches. Rehearsal-based methods maintain a small replay buffer by storing a limited number of representative exemplars from previous tasks [lopez2017gradient, rolnick2019experience, de2021rep]. However, the number of stored exemplars is typically far smaller than that of samples in the current task, creating a severe data imbalance that biases learning toward new classes and thus exacerbates catastrophic forgetting. Restriction-based methods [nguyen2017variational, ahn2019uncertainty] aim to mitigate catastrophic forgetting by constraining model updates during new task learning. A representative example is knowledge distillation [rebuffi2017icarl, simon2021learning], which preserves previous knowledge by aligning the current model’s output with those of its earlier versions. Such constraints introduce a trade-off: the model must balance adherence to past knowledge with adaptation to new data. This stability-plasticity dilemma fundamentally limits the achievable performance of restriction-based approaches.
Figure 1: Parameter-accuracy comparison of different CIL methods on ImageNet-100 with ResNet-18. DLC achieves the best balance between parameter efficiency and accuracy compared to prior SOTA methods. iCaRL-DLC, WA-DLC, and BiC-DLC mean integrating DLC into the corresponding methods, respectively.
To address this, expansion-based CIL methods [mallya2018piggyback, serra2018overcoming, xu2018reinforced] allocate task-specific architectural blocks, creating isolated parameter spaces for new knowledge and freezing previously learned parameters. While this isolates interference between tasks, the scalability limits as each new task requires its own feature extractor, causing parameters to grow linearly and incurring heavy storage costs. Recent pre-trained models-based methods [zhou2024continual] attempt to reduce this cost through lightweight fine-tuning modules. Yet, most PTM-based designs rely heavily on vision transformers with massive parameters (86M), and their fine-tuning modules can even surpass the size of conventional models. For instance, the introduction of a single Adapter module [zhou2024revisiting] incurs an parameter overhead (15M) greater than that of a complete ResNet-18 (11M). In addition, these PTM-based expansion methods face practical limitations, as PTMs are not universally available in practical applications. Therefore, expansion-based methods raise concerns despite strong performance. Their long-term sustainability and scalability in lifelong learning are questionable. This exposes a new quandary to CIL: How can we maintain high accuracy without a prohibitive linear increase in parameters as the number of tasks grows?
To address this challenge, we propose the Deployment of extra LoRA Components (DLC). Rather than expanding the feature extractor model, DLC adopts a plug-and-play extension paradigm. Specifically, we attach task-specific Low-Rank Adaptation (LoRA) plugins to the deep layers of the base model. These plugins are updated exclusively during their corresponding task’s training phase, which occurs after the base model has been trained. This design ensures that the DLC plugins are trained independently, thus enabling their seamless and non-interfering integration into other methods. During inference, the base model loads all LoRA plugins to extract their feature representations, and they are aggregated into a composite one for final classification. However, non-target LoRA plugins that were not adequately updated on the dataset corresponding to the sample’s task may introduce noise or less informative residuals into the composite representation. So we propose a dedicated lightweight weighting unit before the classifier. This unit automatically assigns importance weights to different LoRA plugin outputs. DLC could enhance many non-expansion CIL approaches, such as knowledge distillation or replay methods. Experimental results on various benchmark datasets show that our method consistently improves the performance of replay- and distillation-based baselines. For instance, it could achieve up to 8% accuracy gains on ImageNet-100 using ResNet-18 and on CIFAR-100 using ResNet-32, while only increasing the total number of parameters by 4% and 11%, respectively. Furthermore, under the same memory budget, DLC outperforms existing state-of-the-art baselines. Our contributions are summarized as follows:
  • We propose a plugin extension framework DLC, which can be seamlessly integrated into diverse CIL methods without architectural modification by injecting task-specific residuals into deep layers via LoRA.
  • We introduce a lightweight weighting unit for inference, which adaptively suppresses irrelevant LoRA activations according to their learned importance. This adaptive filtering mechanism effectively mitigates feature interference between old and new tasks, leading to more stable and accurate incremental classification.
  • Extensive experiments show that DLC significantly boosts various CIL methods’ performance. Meanwhile, under the fixed memory budget, the DLC-enhanced distillation method could surpass SOTA baselines.
2 Related Work
DLC can be viewed as a specialized form of expansion-based CIL that augments base methods with task-specific plugins. This section first introduces methods facing the stability-plasticity dilemma that are compatible with DLC, followed by a review of expansion approaches.
2.1 CIL with stability-plasticity dilemma
To mitigate catastrophic forgetting, many methods implement constrained parameter updates during model optimization [nguyen2017variational, bian2024make, cui2025few]. A prominent strategy in this category is knowledge distillation [simon2021learning, huang2024etag]. It protects the previously acquired knowledge through a modified distillation loss from the old to the new model. iCaRL [rebuffi2017icarl] mitigates forgetting by distilling the logits from the old model into the new model. Subsequent methods like BiC [wu2019large] and WA [zhao2020maintaining] further propose bias correction and weight alignment techniques to calibrate the classifier. Moreover, these methods preserve extra knowledge of prior classes by employing replay buffers [shaul2024teal, aghasanli2025prototype] to save exemplars in training. Furthermore, recent research has developed plug-and-play auxiliary loss functions [gao2025maintaining] that can be readily integrated into existing methods to enhance their accuracy. However, the fixed capacity of the feature extractor poses a fundamental constraint. The model is inherently limited in its ability to retain knowledge from past tasks and learn new tasks effectively, forcing it to seek a balance between stability and plasticity through a trade-off between classification and distillation/regularization losses, which ultimately caps its achievable accuracy.
2.2 Expansion-Based CIL
Such methods expand specialized parameters for model [mallya2018piggyback, serra2018overcoming, xu2018reinforced]. DER and TagFex [yan2021dynamically, zheng2025task] train an independent representation model on the new task for feature expansion, while the models of old tasks are frozen to retain the learned knowledge. MEMO [zhou2022model] splits the feature extractor into a generic shallow module and several task-specific deep modules. By increasing task-specific modules with fewer parameters, MEMO could save more memory space for storing replay samples than model-level expansion methods. Expansion-based methods achieve notably higher accuracy ceilings compared to fixed-capacity models. However, these methods result in a linear increase in the total number of parameters with tasks, imposing a significant storage overhead. Specifically, in the case of TagFex, the introduced attention module further increases the parameter overhead of the expansion. Even MEMO still expands over 70% of the extractor’s parameters for every new task. Emerging research leverages pre-trained models (PTMs) [zhou2024revisiting, mcdonnell2024ranpac, tan2024semantically] for CIL, mitigating catastrophic forgetting by leveraging knowledge encapsulated in PTMs. This has spurred the development of specialized parameter expansion methods designed for PTMs, which involve training lightweight fine-tuning modules like adapters [zhou2024expandable, sun2025mos]. However, such methods paradoxically result in higher overall parameter counts rather than a reduction. This is primarily due to the large parameter size of ViT backbones—significantly exceeding that of ResNet—combined with the non-negligible overhead of their tuning modules. An adapter introduces parameters equivalent to 15% of the base ViT’s total. Furthermore, PTM is not available in many CIL cases. These parameter-efficient module expansion methods depend on the rich prior knowledge embedded in PTMs and cannot be directly transferred to non-pre-trained models.
Figure 2: The proposed DLC framework. Left: Training. When a new task $t$ arrives, a dedicated plugin set $L_{t}$ is created. Then sequentially train the feature extractor and $L_{t}$. Right: Test. The feature extractor sequentially loads all task plugins $L_{t}$ to produce enhanced representations, which are concatenated, weighted by a gating unit, and classified.
3 Preliminaries
3.1 Problem Formulation of CIL
A CIL model aims to learn new classes while retaining previous knowledge from the continuously evolving data streams $D_{1},D_{2},\cdots,D_{t}$ [de2021continual], where $D_{i}=\{(\mathbf{x}_{j},y_{j})\}_{j=1}^{n_{i}}$ containing $n_{i}$ instances. One class will appear only once on a given $D_{i}$ and will not be repeated. At each incremental learning stage $t$, only the current data stream $D_{t}$ and a few exemplars from the previous classes [rebuffi2017icarl] can be used for model training. For any stage $t$, the model has learned $D=D_{1}\cup D_{2}\cup\cdots\cup D_{t}$ streams, We aim to find a model $f(\mathbf{x}):\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes empirical risk across all learned classes, where $\mathcal{X}$ and $\mathcal{Y}=Y_{1}\cup Y_{2}\cdots\cup Y_{t}$ represent the sample’s feature and class spaces, respectively. The effectiveness of the model is evaluated by the test dataset $D^{test}=D_{1}^{test}\cup D_{2}^{test}\cup\cdots\cup D_{t}^{test}$, which can be expressed in Eq.(1):
$f^{*}(\mathbf{x})=\arg\min_{f\in\mathbb{H}}\mathop{\mathbb{E}}\nolimits_{(\mathbf{x},y)\in D^{test}}[\mathbb{I}(f(\mathbf{x})\neq y)],$ (1)
where $\mathbb{H}$ is the hypothesis space of model and $\mathbb{I}(\cdot)$ denotes the indicator function.
3.2 –
Consistent Knowledge Preserved in CIL Considering the classic CIL scenario with only a single feature extractor. The objective function used in non-CIL settings is the cross-entropy loss $\mathcal{L}_{\text{CE}}$, while restriction-based CIL methods incorporate an additional restriction term $\mathcal{L}_{\text{MEM}}$ that enforces the preservation of knowledge acquired from previous tasks. This composite objective function can be formally expressed as:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{MEM}}.$ (2)
As a canonical example of restriction-based methods, we examine logit-based distillation, which ensures consistency between the logit distributions produced by the old and new models for known classes. Specifically, the restriction term $\mathcal{L}_{\text{MEM}}$ is defined as a distillation loss $\mathcal{L}_{\text{KD}}$. Let $\hat{q}(\boldsymbol{x})$ and $q(\boldsymbol{x})$ represent the logits of the teacher and student models. The cross-entropy version of $\mathcal{L}_{\text{KD}}$ is defined as:
$\mathcal{L}_{\text{KD}}=-\sum_{i}\hat{q}_{i}(\boldsymbol{x})\log(q_{i}(\boldsymbol{x})),$ (3)
and the KL divergence version is defined as:
$\mathcal{L}_{\text{KD}}=\tau^{2}\sum_{i}\hat{q}_{i}(\boldsymbol{x})\log(\frac{\hat{q}_{i}(\boldsymbol{x})}{q_{i}(\boldsymbol{x})}),$ (4)
where the $\tau$ is the temperature scalar.
Moreover, most methods are complemented by a replay buffer that stores a subset of representative exemplars from previous tasks. These exemplars are interleaved with data from the current task during training to help maintain performance on past tasks. Specifically, when learning the $t$-th task, the model is conventionally trained only on the new dataset $D_{t}$. With the replay buffer, the training dataset is supplemented with $D_{\text{exp}}$ retrieved from the buffer, forming a combined training set $D_{\text{train}}=D_{t}\cup D_{\text{exp}}$.
Although the model’s parameters are dynamically updated throughout the incremental learning process, the use of knowledge distillation and the replay buffer ensures that a portion of knowledge from previous tasks remains preserved. Replay ensures the model continues to minimize the empirical risk on a representative subset of past tasks, anchoring the parameters in regions consistent with previous data. Concurrently, knowledge distillation functions as a functional regularizer, ensuring that the mappings for old classes remain stable by aligning the output logits (or distributions) with those of the prior model. Consequently, a model trained with the replay buffer and distillation retains significant knowledge about past classes, functioning as a capable but suboptimal ”generalist.” We posit that the knowledge stored in the model can be efficiently activated and tuned with minor task-specific parameters.
4 The Proposed Method
While the use of rehearsal and restriction enable the model to retain knowledge from previous tasks, an additional objective or imbalanced exemplars could create an optimization trade-off. This results in a suboptimal convergence point for the cross-entropy loss or past data, ultimately capping the model’s classification performance.
DLC proposes a distinctive paradigm, as shown in Fig. 2. We regard the model trained with a replay buffer and distillation as a foundation model, which stored knowledge across all tasks (albeit with suboptimal accuracy). Then refine it through residuals from each task-specific LoRA plugin to enhance accuracy. The details of the proposed DLC framework are elaborated in the following subsections.
4.1 LoRA for Convolutional Layers
A basic CIL model typically comprises a feature extractor $\phi(\boldsymbol{x})$ followed by a classifier $\boldsymbol{W}$. Most model expansion methods focus on expanding the $\phi(\boldsymbol{x})$ to achieve feature expansion. To maximize parameter efficiency, our approach avoids expanding the entire $\phi(\boldsymbol{x})$ or its deeper layers. Instead, we append parameter-efficient plugins after selected network layers within the $\phi(\boldsymbol{x})$ to adapt the representations produced by the base model. We specifically adopt Low-Rank Adaptation (LoRA) [hu2022lora] due to its minimal parameter size and its architectural flexibility.
The standard LoRA method was originally designed for linear layers. Given that most non-pre-trained CIL models are based on Convolutional Neural Networks (CNNs), it is necessary to change the original LoRA formulation to the convolutional operations fundamental to CNNs. For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation. By analogy to LoRA, ConvLoRA constructs a low-rank adaptation for a convolutional weight tensor via a sequence of two convolutional operations. The forward pass first processes the input with $\mathbf{A}$:
$\boldsymbol{z}=\mathbf{A}\ast\boldsymbol{x},$ (5)
where $\mathbf{A}\in\mathbb{R}^{r\times C_{\text{in}}\times K\times K}$ projects the feature to a low-rank space. Then followed by kernel B:
$\boldsymbol{h}_{\text{lora}}=\mathbf{B}\ast\boldsymbol{z},$ (6)
where $B\in\mathbb{R}^{C_{\text{out}}\times r\times 1\times 1}$ restores channel dimensionality. In summary, the forward pass with ConvLoRA is:
$\boldsymbol{h}^{\prime}=W_{\text{conv}}\ast\boldsymbol{x}+\frac{\alpha}{r}(B\ast(A\ast\boldsymbol{x})),$ (7)
where $\alpha$ is a scaling hyperparameter, $r$ is the low-rank.
4.2 Deploy LoRA as Task-Specific Plugins
LoRA cannot be directly applied in non-PTM settings, not only because PTMs provide diverse knowledge. Without a frozen PTM backbone, the output of the model changes with parameter updates, leading to unstable layer inputs for LoRA. However, LoRA becomes applicable in replay- and distillation-based methods, as these approaches effectively constrain the variation of the model’s outputs on old tasks, thereby ensuring that the input features to LoRA remain close to the distribution encountered during its training phase. Specifically, after training on task $t$, the output deviation of the model at any layer $\ell$ will not exceed $K_{\ell}\Gamma_{t}$ (The detailed proof is provided in the supplementary material). Here, $K_{\ell}$ is determined by hyperparameters such as the network structure and distillation temperature, while $\Gamma_{t}$ characterizes the intensity of distribution discrepancy induced by selected distillation and replay strategies. This upper bound limits the maximum feature drift magnitude of the model on the old tasks’ distribution.
Similar to expansion-based methods, our approach incorporates independent LoRA plugins for each task. Specifically, for each task $t$, we create a dedicated set of $k$ LoRA plugins for $\phi(\boldsymbol{x})$, denoted collectively as $L_{t}$. The $\phi(\boldsymbol{x})$ equipped with $L_{t}$ is denoted by $\phi^{L_{t}}(\boldsymbol{x})$. In DLC, LoRA plugins are deployed in the final layers of $\phi(\boldsymbol{x})$. Taking a CNN-based architecture as an example, the first LoRA plugin in $L_{t}$ is applied to the last convolutional layer of the $\phi(\boldsymbol{x})$. The remaining LoRA plugins are then sequentially deployed in a backward manner toward earlier layers. This design is motivated by two key considerations. First, features extracted from deeper layers are more semantically meaningful and are directly utilized for final classification decisions. Injecting task-specific residual adaptations via LoRA at this level allows for more direct enhancement of inter-class discriminability. Second, it demonstrates better compatibility with distillation-based objectives. When applied to shallow layers, the residual modifications introduced by LoRA would cause alterations in low-level features that propagate through the entire network, consequently amplifying the divergence in the final output logits. This increased discrepancy substantially raises the optimization difficulty for distillation losses, which rely on logit alignment between old and new models.
In DLC, the LoRA plugins and $\phi(\boldsymbol{x})$ undergo a decoupled, two-phase training procedure. Specifically, upon the arrival of a new task $t$, we first train only the $\phi(\boldsymbol{x})$ and classifier $\boldsymbol{W}$, keeping all LoRA plugins frozen, to facilitate adaptation to the new task distribution. After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and the $L_{t}$ are trained with the same data. Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge from forgetting. The training objective for all LoRA plugins incorporates the cross-entropy loss and the auxiliary loss [yan2021dynamically, zhou2022model] commonly employed in expansion-based methods, formulated as follows:
$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}.$ (8)
The two-phase training procedure ensures that the LoRA plugins are trained without interference from distillation losses, enabling specialized learning for the current task and thereby improving its accuracy. Since the $\phi(\boldsymbol{x})$ remain frozen during LoRA updates, the training process of LoRA plugins exhibits stability. Furthermore, it prevents the $\phi(\boldsymbol{x})$’s training from being affected by LoRA, thereby maintaining the plug-and-play characteristic of DLC.
During the inference phase, all LoRA plugins are activated to extract their respective feature representations. These representations are then concatenated to form an integrated feature vector, which enhances the base model through a multi-task feature ensemble. For a linear classifier $\boldsymbol{W}$, the classification process can be formulated as:
$\hat{y}=\boldsymbol{W}[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),...,\phi^{L_{T}}(\boldsymbol{x})]$ (9)
where $T$ denotes the maximum index of known tasks. Compared to the model expansion, LoRA plugin expansion can bring significant parameter advantages. Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while a ConvLoRA only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$. For expanding one layer of CNN over $T$ tasks, considering that in a CNN in CIL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7). Since $r$ is low-rank, it can also be regarded as a constant (generally set to 8 or 16). Then, using LoRA expansion reduces the parameter space complexity of expanding one layer of CNN from $\Theta(TC^{2})$ to $\Theta(TC)$. Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can bring a huge reduction in parameters.
4.3 Dynamic Weighting Unit
For a sample from task $t$, the LoRA plugins trained on other tasks is considered non-target plugins. Given that the replay buffer contains only a minimal fraction of the original data for each task, these non-target plugins are not trained on the full data from task $t$. In particular, plugins trained on tasks 1 to t1t-1 never encounter samples from the task $t$. Consequently, the residuals extracted by these plugins could be largely irrelevant to the task $t$, thus introducing uninformative features or even detrimental noise to the classifier.
To mitigate the influence of uninformative residuals from non-target plugins, we will create a weighting unit corresponding to concentrated representations when training on a new task. This unit comprises two linear layers followed by a sigmoid function $\sigma(\cdot)$, denoted as $\omega(\mathbf{x})=\sigma\Big(\mathbf{W}_{2}\text{ReLU}\big(\mathbf{W}_{1}\mathbf{x}\big)\Big),\mathbf{x}\in\mathbb{R}^{k},\mathbf{W}_{1}\in\mathbb{R}^{k\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times k}$, where the sigmoid function normalizes the output to the [0, 1] range, indicating their importance weights. The concatenated representations are passed through this unit before being fed into the classifier, then weighted representations is derived from $\tilde{\mathbf{h}}^{(i)}=\mathbf{\omega}^{(i)}\odot\mathbf{h}^{(i)}$, where $\odot$ is element-wise multiplication operator. The weighting unit can be regarded as a gate for the classifier. Jointly trained with the classifier, it helps the model adapt to the high-dimensional feature space and automatically assigns lower weights to uninformative or noisy features.
The automatically trained weighting unit fails to explicitly model the varying importance of representations from different tasks. For a sample $i$ in task $t$, we denote the representations adapted by LoRA plugins trained on tasks 1 to t1t-1 as $\mathbf{h}^{(i)}_{\text{pre}}$ and those adapted by the remaining LoRA plugins as $\mathbf{h}^{(i)}_{\text{pos}}$, respectively. Then $\mathbf{h}^{(i)}_{\text{pre}}$ should be assigned lower weights as corresponding LoRA plugins were not optimized on task $t$. However, since the ground-truth task index of each sample is unavailable during the CIL testing, we cannot explicitly control the weighting unit based on task identity. Instead, we employ a supervised regression training to guide the weighting unit to approximate the desired weighting scheme. We decompose the weight $\mathbf{\omega}^{(i)}$ for sample $i$ into $\mathbf{\omega}^{(i)}_{\text{pre}},\mathbf{\omega}^{(i)}_{\text{pos}}$] Then the decision process of the linear classifier can be rewritten as follows:
$\boldsymbol{q}^{i}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})+\mathbf{W}_{\text{pos}}(\mathbf{\omega}^{(i)}_{\text{pos}}\odot\mathbf{h}^{(i)}_{\text{pos}})$. (10)
Formally, the contribution of $\mathbf{h}^{(i)}_{\text{pre}}$ to the classifier’s logits can be expressed as $\mathbf{\xi}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})$. To generate effective weights, an importance-aware optimization objective could be incorporated into the training process, denoted as $\mathcal{L}_{\text{IA}}$. This loss function employs the Mean Squared Error (MSE), expressed as follows:
$\mathcal{L}_{\text{IA}}=\frac{1}{k}{\|\mathbf{\omega}-\mathbf{\omega}_{\text{ideal}}\|}_{2},$ (11)
where the $\mathbf{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\mathbf{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\mathbf{\omega}^{(i)}_{\text{ideal-pre}}=0$. In the ideal training case, the model has $\mathbf{\omega}^{(i)}_{\text{pos}}\to 1$ and $\mathbf{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits. Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed. After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the LoRA training, and $\mathcal{L}_{\text{LoRA}}$ in Eq. (8) is extended as follows:
$\mathcal{L}_{\text{LoRA}}=\mathcal{L}_{\text{CE}}+\mathcal{L}_{\text{aux}}+\mathcal{L}_{\text{IA}}.$ (12)
Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison. ‘-DLC’ indicates the method using the DLC framework. ‘LoRA #P’ denotes the total parameter count of the LoRA plugins, and ‘Ext #P’ refers to the total parameter count of the feature extractor.
Methods CIFAR-100 Tiny-ImageNet ImageNet-100
B5 Inc5 B10 Inc10 B50 Inc10 B40 Inc40 B50 Inc10
$\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $_{T} A¯\overline{A}$ $A_{T}$
LWF 34.04 14.03 43.10 23.09 38.83 22.56 45.23 32.32 44.84 30.22
LWF-DLC 34.20 15.29 43.38 23.30 39.20 22.95 47.86 33.98 46.60 32.10
Replay 58.28 38.20 56.82 39.31 51.08 40.88 51.46 37.79 55.22 43.16
Replay-DLC 58.76 38.69 59.24 41.81 53.72 43.61 51.65 38.66 59.76 49.62
iCaRL 59.40 39.94 59.67 41.29 55.94 44.44 52.32 36.22 62.72 53.55
iCaRL-DLC 59.54 41.63 60.06 44.31 56.61 46.35 53.28 39.20 65.03 54.06
BiC 56.45 32.77 61.14 42.58 53.45 36.04 54.12 39.68 66.02 49.82
BiC-DLC 58.27 41.07 61.87 43.65 54.46 46.23 54.86 43.09 71.53 57.24
WA 62.72 45.96 65.82 51.80 65.04 56.94 53.68 39.69 65.53 56.55
WA-DLC 62.84 46.22 66.79 52.70 65.43 57.71 55.31 43.33 73.84 64.40
LoRA #P 102400 51200 30720 409600 491520
Ext #P 464154 464154 464154 11168832 11168832
5 Experiments
This section conducts extensive experiments. Sec. 5.2 evaluates the validity of DLC on replay and distillation-based baselines. Sec. 5.3 evaluates the compatibility of DLC on other data augmentation and plug-and-play loss function for CIL. Sec. 5.4 investigates the performance of baselines under memory-aligned case. An ablation study is conducted in 5.5 to analyze the contribution of the weighting unit.
5.1 Experimental Setup
Datasets and Split. We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [krizhevsky2009learning], ImageNet-100 [deng2009imagenet], and Tiny-ImageNet-200 [vinyals2016matching], where the number following the dataset is the total number of classes included. Following the ‘B/Base-m, Inc-n’ rule proposed by [de2021continual], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 INC5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage. For a fair comparison, we ensure that the training and test datasets are identical for all methods. Detailed information about the dataset is placed in the supplementary materials.
Implementation Details. All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [zhou2023pycil] and Gao et al. [gao2025maintaining]. We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets. The optimizer employs SGD with a fixed learning rate of 0.1, incorporating a multi-step schedule. The batch size is configured at 128 for CIFAR-100 and 256 for all other datasets. The number of training epochs was set consistent with the baseline method. All the methods are run using the same random number seed.
Evaluation Metrics. Following the benchmark protocol settings [rebuffi2017icarl], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks. We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
5.2 Evaluation on Non-Expansion Baselines
In this section, we evaluate the performance improvement brought by the proposed DLC framework when integrated with established distillation-based baselines. We select four representative methods—LwF [li2017learning], iCaRL [rebuffi2017icarl], WA [zhao2020maintaining], and BiC [wu2019large] —where LwF (without a replay buffer) is used to assess DLC’s enhancement effect in a pure distillation setting. Additionally, a simple Replay method is included to examine DLC’s contribution in scenarios relying solely on the herding rehearsal without distillation. We report the $\overline{A}$ and $A_{T}$ of these established methods with and without DLC across three datasets under varying task configurations. In all experiments, the replay buffer size is fixed at 2,000 exemplars. DLC expands only one LoRA plugin per task, applied to the last convolutional layer of the feature extractor. We also report the total parameters required by the LoRA plugins and the size of the entire feature extractor for comparison. The results are summarized in Table 1.
The results show that DLC significantly improves all classical replay or distillation-based methods with only one LoRA plugin per task. Moreover, the parameter overhead introduced by DLC is negligible. DLC achieves up to 8% accuracy improvement while expanding only 4.4% to 22% parameters equivalent to the feature extractor, highlighting its practical value in CIL. Fig. 3 shows the confusion matrix heatmap of iCaRL and iCaRL-DLC as an example. DLC increases the thermal intensity along the diagonal for most classes, demonstrating that it could enhance the model’s accuracy across most tasks.
Figure 3: Confusion matrix heatmaps of iCaRL with and w/o DLC on CIFAR-100 B10 Inc10.
5.3 Compatibility with other Plug-and-Play Enhancements
Figure 4: Comparison of $A_{T}$ with/without DLC on CIFAR-100 when enhanced with AutoAugment (AA) and $\mathcal{L}_{\text{inter-intra}}$.
In addition to DLC, other plug-and-play strategies have been proposed to enhance CIL methods. For example, the data augmentation technique AutoAugment [cubuk2018autoaugment] has been employed in the CIFAR-100 dataset by [wang2022foster, zheng2025task]. Meanwhile, two additional training losses $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ specifically designed for logit-based distillation methods have been introduced in [gao2025maintaining]. This subsection evaluates the compatibility of DLC with such enhance strategies. Specifically, we compare the accuracy of methods using AutoAugment and the loss $\mathcal{L}_{\text{inter-intra}}$ with and without DLC. As shown in the Fig. 4, DLC consistently brings significant improvements, demonstrating its compatibility with other plug-and-play enhancements.
5.4 Comparison under Fixed Memory Budget
Prior research [zhou2024class] has demonstrated that model-level expansion-based methods achieve state-of-the-art performance across all benchmarks. However, these approaches overlook the substantial memory overhead induced by dynamic architecture growth. Such methods implicitly introduce additional memory budgets, i.e., a model buffer for retaining parameters of the expanded model, which grants them a considerable and unaccounted advantage over other methods that do not store expansion models or adopt more lightweight expansion strategies.
Following the protocol established in [zhou2022model], this section presents a memory-aligned comparison of CIL baselines without pre-training. Under a fixed 23.5 MB memory budget, the memory not consumed by a method’s parameters is allocated to replay exemplars. To balance this, methods with fewer parameters store more exemplars, and vice versa. This ensures an equitable evaluation. In addition to the baselines used in Section 5.2, we add several state-of-the-art approaches for comprehensive comparison, including FOSTER [wang2022foster], DER [yan2021dynamically], MEMO [zhou2022model], and TagFex [zheng2025task]. Among them, DER, MEMO, and TagFex have incorporated the weight alignment from WA in their implementations. As summarized in Table 2, deploying DLC to the distillation-based method WA under the memory-aligned setting yields the highest classification accuracy.
Table 2: Performance comparison on CIFAR-100 with aligned memory cost. ‘#P’ represents the number of parameters (million). $\varepsilon$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB). All methods use ResNet-32 as the backbone.
Method CIFAR100 B10 Inc10
MS #P $\varepsilon$ $\overline{A}$ $A_{T}$
iCaRL 23.5 0.46 7431 70.74 58.12
WA 23.5 0.46 7431 69.25 59.00
BiC 23.5 0.46 7431 70.29 59.20
FOSTER 23.5 0.46 7431 72.18 59.29
DER+WA 23.5 4.60 2000 71.24 60.26
MEMO+WA 23.5 3.62 3312 72.32 61.92
TagFex+WA 23.5 5.62 700 74.00 64.10
WA-DLC 23.5 0.72 7000 75.02 64.79
5.5 Ablation Study on the Weighting Unit
This subsection evaluates the impact of the proposed weighting unit on model performance. We compare the final accuracy of DLC-enhanced iCaRL, WA, and BiC with and without the weighting unit, as shown in Fig. 5. Results indicate that the proposed weighting unit contributes a certain accuracy gain. Although this improvement is relatively modest compared to the gains achieved by the LoRA plugin, the weighting unit could effectively direct the model’s attention to representations adapted by the pivotal LoRA plugin or key features of the sample, thereby enhancing their discriminative power. Specifically, we visualize the activation patterns of the last convolutional layer with the target task LoRA plugin using Grad-CAM [selvaraju2017grad], as illustrated in Fig. 6. For a “road” sample, despite both models correctly classifying it, the model equipped with the weighting unit effectively focuses on the road surface. Regarding a “pear” sample, when the weighting unit is ablated, the activation patterns remain inactive, indicating that representations adapted by the target LoRA contribute negligibly to classification. In contrast, with the weighting unit, the model not only correctly utilizes residuals from the target plugin but also focuses on discriminative regions, such as the cross‑section of the pear to improve the accuracy.
Figure 5: Comparison of $A_{T}$ with and without the weighting unit for DLC-enhanced methods on CIFAR-100.
Figure 6: Grad-CAM visualizations for the last convolutional layer integrated with the target-task LoRA plugin. The model is iCaRL-DLC trained on the CIFAR-100 B10 Inc10. ‘Original’ denotes the input image, while ‘W/O weight’ and ‘Weighting’ represent the results without and with the weighting unit.
6 Conclusion
While many class incremental learning methods avoid parameter-intensive expansion by relying on replay or distillation, they often face inherent performance limitations due to imbalanced replay buffers or constraints imposed by distillation. Our work presents DLC, a plug-and-play expansion method that enhances the performance of the base model by task-specific LoRA plugins. These plugins are trained separately and frozen to prevent interference from distillation loss and protect learned knowledge. A trainable weighting unit further suppresses irrelevant features from non-target plugins. Experimental results show that DLC achieves consistent enhancements in non-expansion CIL methods with a few parameter overhead. Future improvements could involve designing a more effective weighting unit to better integrate features or deploying LoRA in earlier layers for enhanced adaptability.